(6a-54)-(5a+27)=23
Answer by askmemath(368)

(6a-54)-(5a+27)=23
When you see a sign(whatever the sign may be) in front of a pair of brackets, you can remove the brackets by mlutiplying each of the inside terms with the sign outside.
Thus
6a-54 - 5a-27 = 23
a-81= 23
a = 81+23 (Adding 81 on both sides)
a =104